2096) Two noninteracting particles 1 and 2, each of mass m, are in a 1-D infinite square well ol width a. If one is in the state Vin and the other in the state (n! /), calculate C(xI-x), assuming (a) (6%) they are distinguishable particles, (b) (7%) Ihey are identical bosons, and (c) (796) Ihey arc identical fermions. 4.
Answer #1

langle(x_1-x_2)^2 angle=langle x_1^2 angle+langle x_2^2 angle-2langle x_1x_2 angle=langle x_1^2 angle+langle x_2^2 angle-2langle x_1 angle langle x_2 angle

because for 2 independent(non-interacting) variables A and B, langle AB angle=langle A anglelangle B angle

Also, single particle wavefunctions are

Vasin (ar) CL

Now, lets begin!

(a) If particles are distinguishable (which is a non-realistic situation for quantum particles, but anyway) there is no condition on the overall wavefunctions, hence we have simply

Psi_{nl}(x_1,x_2)=psi_n(x_1)psi_l(x_2)=rac{2}{a}sinleft ( rac{npi x_1}{a} ight )sinleft ( rac{lpi x_2}{a} ight )

Note: Because we can label the particle(possible for distinguishable), I have explicitly chosen "1" in n-th state and "2" in l-th state.

T12

Some simplification can be done without actually expanding

langle(x_1-x_2)^2 angle=int x_1^2psi_n^2,dx_1+int x_2^2psi_l^2,dx_2-2int x_1psi_n^2,dx_1int x_2psi_l^2,dx_2

3 2n 22 212z?) 3 -22

6 2n2r2 212z?

(b,c) Can't label them anymore.

Psi_{nl}(x_1,x_2)=rac{1}{sqrt2}left [psi_n(x_1)psi_l(x_2)pmpsi_l(x_1)psi_n(x_2) ight ]

+ for Bosons and - for Fermions. On doing the calculation you will get the same answer as (a)